# algorithmic

### v2

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a variant of reinforcement learning (RL) that learns from human feedback instead of relying on an engineered reward function. Building on prior work on the related setting of preference-based reinforcement learning (PbRL), it stands at the intersection of artificial intelligence and human-computer interaction. This positioning offers a promising avenue to enhance the performance and adaptability of intelligent systems while also improving the alignment of their objectives with human values. The training of large language models (LLMs) has impressively demonstrated this potential in recent years, where RLHF played a decisive role in directing the model’s capabilities toward human objectives. This article provides a comprehensive overview of the fundamentals of RLHF, exploring the intricate dynamics between RL agents and human input. While recent focus has been on RLHF for LLMs, our survey adopts a broader perspective, examining the diverse applications and wide-ranging impact of the technique. We delve into the core principles that underpin RLHF, shedding light on the symbiotic relationship between algorithms and human feedback, and discuss the main research trends in the field. By synthesizing the current landscape of RLHF research, this article aims to provide researchers as well as practitioners with a comprehensive understanding of this rapidly growing field of research.
@@ 1 Introduction
In reinforcement learning (RL), an agent traditionally navigates through an environment and attempts to make optimal decisions (i.e., action choices) through a process of trial and error. Whether a decision is optimal or not is determined solely by reward signals. These signals have to be defined by a system designer based on measurements of the agent’s performance, ensuring that the learning agent receives the necessary feedback to learn the correct behavior. Designing a reward function, however, is challenging. Indeed, success is hard to formally define and measure in many applications. Beyond that, a sparse signal of success may not be well suited for agent learning – resulting in the need for reward shaping (Ng et al., 1999), where the reward signal is transformed into one that is more suitable for learning. This often makes the reward signal more susceptible to spurious correlations, however – behaviors that are rewarded because they are usually correlated with the true objective but are not valuable in themselves. This ultimately cumulates in the issue of reward hacking (Skalse et al., 2022), where learning agents exploit reward-specific loopholes to achieve undesired outcomes while still generating high rewards.
In response to these challenges, reinforcement learning from human feedback (RLHF) has emerged as a practically meaningful alternative that introduces a critical human-in-the-loop component to the standard RL learning paradigm. In a nutshell, RLHF differs from RL in that the objective is defined and iteratively refined by the human in the loop instead of being specified ahead of time. This approach not only has the potential to overcome the limitations and issues of classical RL methods but also has potential benefits for agent alignment, where the agent’s learning goals are more closely aligned with human values, promoting ethically sound and socially responsible AI systems.
@@
RLHF has seen a number of successful applications, advances in methodology, and theoretical insights since the last comparable survey (Wirth et al., 2017). The applications span various domains, including large language model (LLM) fine-tuning (OpenAI, 2022), image generation (Lee et al., 2023), continuous control (Christiano et al., 2017), games (Ibarz et al., 2018), and robotics (Hejna & Sadigh, 2022). At the same time, there have been a lot of developments in the methodology since the last comparable survey (Wirth et al., 2017). Examples of methodological developments include fusing multiple feedback types to leverage their relative strengths (see Section 3.5), enhancing query efficiency through active learning and active query synthesis (see Section 4.1.1), incorporating psychological insights to improve the quality of human feedback (see Section 4.2.1), using techniques such as meta-learning to quickly adapt learned preferences to new tasks using prior data (see Section 5.5.1), and using available preference data more efficiently through approaches such as data augmentation and semi-supervised learning (see Section 5.5.2). Finally, there have been some achievements with regard to theoretical results for the field of RLHF (Section 7), providing new insights but also new questions for the fundamental mathematical problems underlying the modeling of the learning scenario in RLHF.
@@
In this survey, we, therefore, discuss the current state of affairs with regard to the ongoing research in RLHF, classify the current approaches as well as concisely describe their main characteristics, and give a brief overview of the application areas. In the remainder of this section, we will start by discussing the motivation (Section 1.1) and origins (Section 1.2) of RLHF as well as the scope of this survey (Section 1.3) and conclude by outlining the contents of the following sections (Section 1.5).
1.1 Why Human Feedback
@@
In conventional RL, the agent’s objective is defined by a reward function that it aims to maximize (Sutton & Barto, 2018). Specifying this reward function can be challenging, particularly in complex domains: What would be a suitable reward function for a robot assisting humans in a household environment or for autonomous vehicles navigating through a busy urban environment? Moreover, even reward functions that initially seem well-defined can lead to unexpected behaviors due to distributional shifts or over-optimization, raising practical and safety concerns. Learning the agent’s objective from human feedback circumvents reward engineering challenges and fosters robust training, with the reward function dynamically refined and adjusted to distributional shifts as the agent learns.
@@
The field of inverse RL aims to infer reward functions from human demonstrations (Arora & Doshi, 2021). While this can partially resolve reward engineering challenges, it faces inherent difficulties: (i) It is generally not possible to robustly identify rewards from demonstrations (Cao et al., 2021a; Mindermann & Armstrong, 2018), (ii) it is only applicable in scenarios where good demonstrations can be obtained, (iii) it struggles to outperform the demonstrator, and (iv) humans often do not demonstrate the behavior they would prefer a machine to adopt (Basu et al., 2017). Interactive feedback, by contrast, can use active queries to differentiate between human preferences and irrelevant noise, is much easier to provide than demonstrations, does not require near-optimal performance from the human evaluators, and elicits preferences on the behavior that a human would prefer from the machine. Interactive feedback can also be used to complement demonstrations, in which case it can be used to shape and refine capabilities learned through initial training, like behavioral cloning, thereby preventing overfitting to demonstrated behavior (Abramson et al., 2022).
@@
Reward engineering in RL presents significant challenges, as accurately specifying reward functions is notoriously difficult (Amodei et al., 2016; Knox et al., 2023). These challenges can be mitigated by utilizing human feedback, which enables training agents for tasks that are hard to define manually and helps avoid safety issues arising from misaligned rewards (Skalse et al., 2022). Safety issues related to a misalignment between the agent’s and the human’s objectives are studied as the AI alignment problem (Gabriel, 2020), in particular agent alignment and value alignment (Kirchner et al., 2022).
Excessive optimization for poorly specified rewards often leads to unintended behaviors. Examples of such behaviors include exploiting flaws in the environment simulation for higher rewards (Lehman et al., 2020; Baker et al., 2020) or engaging in more general reward hacking (Skalse et al., 2022), where the behavior maximizes the specified reward but deviates from the intended objective. This is evident in cases where agents focus on intermediate rewards without achieving the actual goal (Clark & Amodei, 2016) or prematurely exit a game to avoid negative rewards (Saunders et al., 2018). The root of these issues is that the reward function does not properly reflect the actual learning task. While these issues may seem trivial in game-like environments, their implications are far more serious in safety-critical contexts such as healthcare and autonomous driving. In these settings, it is crucial to prevent misaligned reward functions from leading to harmful outcomes, like a care robot causing injury or a self-driving car jeopardizing road safety.
RLHF presents a promising approach to enhance alignment (Leike et al., 2018) by enabling agents to learn from human feedback, which is often more closely aligned with the true objective than manually specified rewards. Nonetheless, the effectiveness of RLHF in resolving these alignment issues is debated (Christiano, 2023). Examples of possible pitfalls raised in this debate are that the agent may be incentivized to manipulate the human teacher to provide feedback that is easier to optimize (Armstrong et al., 2020; Carroll et al., 2023) or that the agent may learn to exploit errors in human judgement (Ngo et al., 2024). We refer the interested reader to the survey by Casper et al. (2023) for a more detailed discussion of these issues. Despite this debate, RLHF represents an important early step towards aligning agents with human values and serves as a foundation to build on to further improve agent alignment.
@@ 1.2 The Origins of RLHF
Learning behavior from human feedback has long been studied as a subfield of RL, but methods and terminology have evolved over time. Early methods focused on learning directly from human rewards (Knox, 2012; Isbell et al., 2001; Knox & Stone, 2009), action advice (Maclin et al., 2005), or action critique (Judah et al., 2010). Notable approaches in this area include TAMER (Knox & Stone, 2009; Warnell et al., 2018), which interprets human feedback as samples of the optimal action-value function, and the later COACH (MacGlashan et al., 2017; Arumugam et al., 2019), which interprets human feedback in a policy-dependent way, i.e., as samples of the advantage function. This survey, however, focuses on more indirect approaches to inferring the objective from human feedback.
Reinforcement learning from human feedback (RLHF) in its modern guise has its origin in the setting of preference-based reinforcement learning (PbRL) as introduced independently by Akrour et al. (2011) and Cheng et al. (2011). The original idea of preference-based reinforcement learning (PbRL) is to infer the objective from qualitative feedback, such as pairwise preferences between behaviors or between actions given states, instead of quantitative feedback in the form of numerical rewards. The term RLHF was coined as an alternative later on (Askell et al., 2021; Ouyang et al., 2022; OpenAI, 2022), though initially referring to the same concept of learning behavior from relative feedback.
@@
Disentangling PbRL and RLHF is challenging due to their overlapping use in the literature. For instance, Christiano et al. (2017) themselves are using the term PbRL, yet are often cited as a seminal reference for RLHF (Daniels-Koch & Freedman, 2022; Ouyang et al., 2022). This indicates the interchangeability of these terms. Practically, RLHF is often associated with reward modeling and deep RL, while PbRL is often linked to direct policy optimization in traditional RL settings. This is underlined by Jeon et al. (2020), who characterize PbRL as limited to direct policy learning from preferences. This is in contrast with other sources, however, who include reward learning within the scope of PbRL (Christiano et al., 2017; Wirth et al., 2017).
Despite the overlapping and sometimes conflicting usage, RLHF is increasingly viewed as a generalization of PbRL. While both involve human feedback to define RL objectives, PbRL primarily focuses on relative feedback, such as binary comparisons and rankings. RLHF not only includes these aspects but also extends to a wider range of feedback types (Metz et al., 2023; Yuan et al., 2024). Table 1 gives an exemplary overview of our interpretation of these terms.
@@
In our work, we adopt the viewpoint that RLHF is a broader category that encompasses various approaches where human feedback is used to define the objective of an RL agent. In this definition, RLHF encompasses both PbRL and SSRL. As the definitions and distinctions between these terms are not universally agreed upon, these distinctions are based on our interpretation of the current predominant usage of these terms in the literature.
@@ 1.3 Scope of the Survey
This section outlines the criteria guiding our selection of approaches in the realm of RLHF. We focus on works that rely on a reward model as the sole source of information about the objective. This reward model should then be learned in an interactive, online, scalable, and asynchronous manner. The following will describe each of these criteria in more detail.
@@
We focus on approaches that learn a reward model from human feedback and then use this model to train a policy. Although it is possible to directly optimize a policy from human feedback (Wirth et al., 2017), thereby performing RLHF without reward learning, this approach was almost not practiced for a long time and has only recently gained renewed interest especially in the domain of language model fine-tuning (see Section 6.3). The decomposition into reward learning and policy training offers many conceptual and practical benefits. Among those benefits are the direct applicability of supervised learning techniques for the reward model and the possibility of evaluating the reward model in isolation. In addition to that, the decomposition naturally leads to a form of semi-supervised learning, enabling the agent to use labeled episodes for reward model training while leveraging unlabelled episodes to refine its behavior and explore the environment.
@@
While there are many approaches that include humans in the RL loop, in this survey, we focus on approaches where human feedback is the only source of truth about the objective. This excludes approaches to reward shaping, feature engineering, and other forms of human guidance that are supplementary to a given objective.
@@
We also put an emphasis on providing feedback in an interactive, online manner. This excludes imitation learning, learning from demonstration, and pure inverse RL. While we do not directly cover inverse RL in this survey, combinations of inverse RL methods with interactive improvements of the reward function are in scope and employed by some of the surveyed methods. See Sections 3.3 and 5.5.1 for a discussion of those approaches.
@@
We focus on works in which the human is included in the loop, but the agent is not blocked by the human’s feedback, and the human does not need to be present continuously. This distinguishes RLHF from more direct methods of incorporating a human into the RL loop, and we believe that this is key for practicality and efficiency.
In addition to these criteria, we mainly focus on works published after 2017 since earlier works are surveyed by Wirth et al. (2017). Nevertheless, some works from this period are revisited from time to time in order to elaborate on certain concepts that are still state of the art or have significantly shaped it. Note that while there has been a recent forus on RLHF for LLMs, LLMs are not the primary focus of this work. Instead, we cover RLHF in a broader context, focusing on control applications as well as discussing the implications of RLHF for fine-tuning generative models such as LLMs.
1.4 Prior Surveys
Based on the criteria mentioned in the previous section, we will first differentiate our survey from other surveys in marginally related subject areas sharing the common theme of human-in-the-loop RL. Then, we will describe the differences between our survey and previous surveys or survey-like articles that exist within the RLHF field.
1.4.1 Human-in-the-Loop RL
Wu et al. (2022) Human-in-the-loop ML
Retzlaff et al. (2024) Human-in-the-loop RL (✓) (✓) (✓)
Najar & Chetouani (2021) RL with human advice (✓) (✓) (✓)
Lin et al. (2020a) Social feedback (✓)
Poole & Lee (2024) RL from brain signals
Cruz & Igarashi (2020) Interactive RL for HCI
Osa et al. (2018) Imitation learning
Arora & Doshi (2021) Inverse RL
Bignold et al. (2021) Assisted RL (✓)
Luketina et al. (2019) Language-informed RL
Zhang et al. (2021) Human guidance
Ji et al. (2023b) AI Alignment
Liu et al. (2023c) LLM applications
Ours RLHF
Human participation in machine learning (ML), particularly in guiding machine learners, is a much-studied scenario. This field, commonly referred to as human-in-the-loop ML, can be further divided into subfields based on various criteria, e.g., the ones detailed in Section 1.3. Prior surveys of these subfields are compiled in Table 2 and briefly summarized in the following.
Human-in-the-Loop
Learning from human feedback falls into the domain of human-in-the-loop ML. Wu et al. (2022) survey human-in-the-loop ML in general. They also cover some applications of RLHF (for LLMs in particular) but do not give a detailed overview. Retzlaff et al. (2024) provide a similar overview over human-in-the-loop RL in particular, focusing on human involvement in RL on a more abstract level than our work and not covering RLHF in detail. Similarly broad in scope, Najar & Chetouani (2021) study the setting of RL with human advice, which they define as ‘teaching signals that can be communicated by the teacher to the learning system without executing the task.’ While this setting subsumes RLHF, the broad generality limits the depth in which their survey can cover RLHF approaches.
Interactive RL
Cruz & Igarashi (2020) survey interactive RL from an human-computer interaction (HCI) viewpoint, exploring various ways humans can influence RL agents, with a particular focus on reward definition based on human feedback, without a predefined environmental reward function. Due to the breadth of their survey, they do not cover many works in this area. The survey by Lin et al. (2020a) centers on interactive RL using human social cues, like gestures and spoken language, but does not cover the reward modeling aspect. Similarly, the study by Poole & Lee (2024) examines RL with direct feedback from human brain signals, such as through brain-computer interfaces, also not focusing on reward modeling.
Demonstrations
Learning from demonstrations, in the form of behavior cloning (Osa et al., 2018) and inverse RL (Arora & Doshi, 2021), shares the goal of RLHF to learn behavior from human input. In contrast to RLHF, however, it requires demonstrations of the desired behavior instead of feedback, and these demonstrations are usually not provided interactively and online. This limits their applications and also their final performance due to the availability of near-optimal demonstrations. Nonetheless, imitation and demonstration can be a useful component of an RLHF system but are not the main focus of this survey. However, we will discuss the intersection between these fields in some parts whenever necessary.
Assisted RL
Bignold et al. (2021) review the field of assisted RL, where an agent may receive external information (for example, from a human) that aids it in action selection. While updates to the reward function are one of the possible effects of advice in this setting (in addition to action selection or modifications of the agent’s internal state), it is usually assumed that an initial reward function is given and the extent of the updates is limited to reward shaping or supplementary reward signals. In contrast to RLHF, the external information does not define the task but only helps the agent in achieving it. Closely related to this, Luketina et al. (2019) survey RL assisted by natural language. In addition to this assistance setting, they also discuss approaches that infer a language-conditioned reward function. However, they discuss this setting rather briefly and use techniques from inverse RL and not RLHF.
Guidance
In their survey on human guidance, Zhang et al. (2021) delve into various aspects related to RLHF. Although they touch on aspects such as reward learning, it is not the primary emphasis of their work. Instead, their main focus lies on exploring more immediate approaches that do not involve the learning of a reward model.
AI Alignment
Applications
1.4.2 PbRL and RLHF
Table 3: An overview of prior RLHF-specific surveys and articles with substantial review components. indicates that the aspect is addressed, (✓) indicates that the aspect is partially addressed, while indicates that the aspect is not covered.
Wirth et al. (2017)
(preference-based RL)
Abdelkareem et al. (2022)
(recent advances of PbRL)
Jeon et al. (2020)
(feedback modelling)
Casper et al. (2023)
(open issues in RLHF)
Fernandes et al. (2023).
(language generation)
Metz et al. (2023)
(feedback types)
Yuan et al. (2024)
(feedback types)
Ours (fundamentals, recent advances, and trends)
There have been previous surveys or survey-like articles that are closely related to RLHF. Table 3 gives a brief overview of how these articles differ from ours, which we will explain in more detail below.
Preference-Based RL
Previous surveys in the domain of RLHF often focus on PbRL, where feedback is limited to binary preferences (see Section 1.2). An illustrative example of this is the survey by Wirth et al. (2017), which is a direct precursor to our work. In contrast to our work, they concentrate on binary preferences for trajectories and primarily survey methods that learn policies without deriving a reward model. Since then, the reward-modeling approach has become dominant in the field, and other approaches have extended RLHF to new feedback types. Abdelkareem et al. (2022) give another more recent literature review of PbRL. While this review focuses on reward modeling and includes some recent work, it is far less comprehensive than our review, as many aspects are only touched upon and partly overlap with those of Wirth et al. (2017).
Feedback Types
Although not a survey per se, Jeon et al. (2020) propose reward-rational implicit choice as a unifying framework to comprehend many previous studies in PbRL and RLHF. To illustrate its generality, they overview different feedback types used in previous work and explain how they fit into their framework. The concurrent works by Metz et al. (2023) and Yuan et al. (2024), which are also not strictly surveys, propose frameworks for studying user interaction and interface design for multiple feedback types. As part of their work, they provide a classification of feedback types and a brief overview of RLHF approaches. Metz et al. (2023) have a stronger focus on the feedback interface and on learning from multiple feedback types simultaneously, discussing properties of feedback types and proposing a standard encoding for them. On the other hand, Yuan et al. (2024) also include an offline RLHF benchmark and have a stronger focus on the reward learning aspect, focusing on the entire learning pipeline. Nevertheless, many facets of RLHF are not dealt with at all in those studies, as they are not primarily survey articles. Our survey has a broader scope and, therefore, has more extensive coverage, going beyond the study of feedback types and discussing more recent work.
Domain-Specific
Fernandes et al. (2023) focuses on human feedback for langauge generation. As a result of their focus, their survey is less comprehensive than this work but discusses some language-specific aspects that do not fall into our scope, such as using feedback models at generation time.
Open Problems
@@ 1.4.2 PbRL and RLHF
All in all, our survey can be seen as the canonical continuation of Wirth et al. (2017), which examines the evolution of the field of PbRL to the more modern and general field of RLHF. This includes a thorough description of the basics as well as an in-depth discussion of current advances and trends in the field.

### v3

# A Survey of Reinforcement Learning from Human Feedback

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a variant of reinforcement learning (RL) that learns from human feedback instead of relying on an engineered reward function. Building on prior work on the related setting of preference-based reinforcement learning (PbRL), it stands at the intersection of artificial intelligence and human-computer interaction. This positioning provides a promising approach to enhance the performance and adaptability of intelligent systems while also improving the alignment of their objectives with human values. The success in training large language models (LLMs) has impressively demonstrated this potential in recent years, where RLHF has played a decisive role in directing the model’s capabilities towards human objectives. This article provides an overview of the fundamentals of RLHF, exploring how RL agents interact with human feedback. While recent focus has been on RLHF for LLMs, our survey covers the technique across multiple domains. We provide our most comprehensive coverage in control and robotics, where many fundamental techniques originate, alongside a dedicated LLM section. We examine the core principles that underpin RLHF, how algorithms and human feedback work together, and the main research trends in the field. Our goal is to give researchers and practitioners a clear understanding of this rapidly growing field.
Contents
@@ 1 Introduction
In reinforcement learning (RL), an agent navigates an environment and learns through trial and error. The optimality of these decisions is determined solely by reward signals, which must be carefully designed based on performance measurements to ensure the agent learns the desired behavior. However, designing effective reward functions presents significant challenges. In real-world applications, success is often difficult to formally define and measure. Moreover, sparse success signals are not well suited for agent learning, often necessitating reward shaping (Ng et al., 1999), where the reward signal is transformed into one that is more suitable for learning. Unfortunately, this transformation can introduce spurious correlations – rewarding behaviors that correlate with desired outcomes without directly achieving the true objective. This can lead to reward hacking (Skalse et al., 2022), where agents exploit such correlations to maximize rewards while producing undesired outcomes.
Reinforcement learning from human feedback (RLHF) has emerged as a practical solution to these challenges by incorporating human feedback directly into the learning process. Unlike traditional RL, where objectives are fixed a priori, RLHF allows humans to define and iteratively refine objectives during training. This approach addresses the limitations of classical RL while promoting better alignment between agent behavior and human values and objectives, enabling the development of more ethically sound, socially responsible, and practically useful AI systems.
@@
RLHF has seen a number of successful applications, methodological advances, and theoretical insights since the last comparable survey (Wirth et al., 2017). The applications span various domains, including large language model (LLM) fine-tuning (OpenAI, 2022), image generation (Lee et al., 2023), music generation (Cideron et al., 2024), continuous control (Christiano et al., 2017), games (Ibarz et al., 2018), and robotics (Hejna & Sadigh, 2022). Alongside these applications, significant methodological advances have emerged. These include fusing multiple feedback types to leverage their relative strengths (Section˜3.5), enhancing query efficiency through active learning and query synthesis (Section˜4.1.1), incorporating psychological insights to improve feedback quality (Section˜4.2.1), using techniques such as meta-learning to quickly adapt learned preferences to new tasks using prior data (Section˜5.5.1), and using available preference data more efficiently through approaches such as data augmentation and semi-supervised learning (Section˜5.5.2). Finally, novel theoretical results (Section˜7) have provided insights as well as new questions about the fundamental mathematical problems underlying RLHF.
While we focus on the core principles underlying RLHF, LLM fine-tuning requires additional considerations due to the unique characteristics of LLMs and their operational scale. We address these considerations without delving deeply into LLM-specific details. Notably, applying RLHF to LLMs is not our primary focus. While RLHF for LLMs remains in scope, we survey RLHF broadly, examining techniques and applications across domains with particular attention to control applications. Nonetheless, many techniques and insights discussed herein apply to LLMs, and we address some unique aspects of LLM fine-tuning. For readers primarily interested in LLM fine-tuning, Appendix˜A provides an LLM-focused introduction that complements this survey and guides the application of the insights presented here to LLM fine-tuning. Note that the literature focused on RLHF for LLMs in particular is vast and rapidly growing, and we cannot provide a comprehensive overview of this area.
@@
This survey therefore provides an overview of current RLHF research, classifies existing approaches, describes their main characteristics, and briefly surveys application areas. In the remainder of this section, we begin by examining the motivation (Section˜1.1) and origins (Section˜1.2) of RLHF, defining the scope of this survey (Section˜1.3), and outlining the structure of subsequent sections (Section˜1.4).
1.1 Why Human Feedback?
@@
In conventional RL, the agent’s objective is defined by a reward function that it aims to maximize (Sutton & Barto, 2018). Specifying this reward function can be challenging, particularly in complex domains: What would be a suitable reward function for a robot assisting humans in a household environment or for autonomous vehicles navigating through a busy urban environment? Moreover, even reward functions that initially seem well-defined can lead to unexpected behaviors due to distributional shifts or over-optimization, raising practical and safety concerns. Learning the agent’s objective from human feedback circumvents reward engineering challenges and fosters robust training, with the reward function being dynamically refined and adjusted to distributional shifts as the agent learns.
@@
The field of inverse RL aims to infer reward functions from human demonstrations (Arora & Doshi, 2021). While this can partially resolve reward engineering challenges, it faces inherent difficulties: (i) it is generally not possible to robustly identify rewards from demonstrations (Cao et al., 2021a; Mindermann & Armstrong, 2018), (ii) it is only applicable in scenarios where good demonstrations can be obtained, (iii) it struggles to outperform the demonstrator, and (iv) humans often do not demonstrate the behavior they would prefer a machine to adopt (Basu et al., 2017). Interactive feedback, by contrast, can use active queries to differentiate between aspects relevant or irrelevant to the human preference, is much easier to provide than demonstrations, does not require near-optimal performance from human evaluators, and elicits preferences about the behavior that a human would prefer from the machine. Interactive feedback also complements demonstrations effectively, refining capabilities learned through initial training methods like behavioral cloning and preventing overfitting to demonstrated behavior (Abramson et al., 2022).
@@
Reward engineering in RL presents significant challenges, as accurately specifying reward functions is notoriously difficult (Amodei et al., 2016; Knox et al., 2023). These challenges can be mitigated by using human feedback, which enables training agents for tasks that are hard to define manually and represents a step towards addressing safety issues arising from misaligned rewards (Skalse et al., 2022). Safety issues related to a misalignment between the agent’s and the human’s objectives are studied as the AI alignment problem (Gabriel, 2020), in particular agent alignment and value alignment (Kirchner et al., 2022).
Excessive optimization of poorly specified rewards frequently produces unintended behaviors. Examples of such behaviors include exploiting flaws in the environment simulation for higher rewards (Lehman et al., 2020; Baker et al., 2020) or engaging in more general reward hacking (Skalse et al., 2022), where the behavior maximizes the specified reward but deviates from the intended objective. This is evident in cases where agents focus on intermediate rewards without achieving the actual goal (Clark & Amodei, 2016) or prematurely exit games to avoid negative rewards (Saunders et al., 2018). The root of these issues is that the reward function does not properly reflect the actual learning task. While such problems appear trivial in game environments, they become critical in safety-critical contexts such as healthcare and autonomous driving, where misaligned rewards could lead to care robots causing injuries or self-driving cars compromising road safety.
RLHF presents a promising approach to enhance alignment (Leike et al., 2018) by enabling agents to learn from human feedback, which is often more closely aligned with the true objective than manually specified rewards. Nonetheless, the effectiveness of RLHF in resolving these alignment issues is debated (Christiano, 2023). Examples of possible pitfalls raised in this debate are that an agent may be incentivized to manipulate the human teacher to provide feedback that is easier to optimize (Armstrong et al., 2020; Carroll et al., 2023) or that the agent may learn to exploit errors in human judgment (Ngo et al., 2024). We refer the interested reader to the survey by Casper et al. (2023) for a more detailed discussion of these issues. Despite these concerns, RLHF represents an important early step towards aligning agents with human values and serves as a foundation to build on to further improve agent alignment.
@@ 1.2 The Origins of RLHF
Learning behavior from human feedback has long been studied as a subfield of RL, but methods and terminology have evolved over time. Early methods focused on learning directly from human rewards (Knox, 2012; Isbell et al., 2001; Knox & Stone, 2009), from action advice (Maclin et al., 2005), or from action critique (Judah et al., 2010). Notable examples include TAMER (Knox & Stone, 2009; Warnell et al., 2018), which interprets human feedback as samples of the optimal action-value function, and COACH (MacGlashan et al., 2017; Arumugam et al., 2019), which interprets human feedback in a policy-dependent way, i.e., as samples of the advantage function. This survey, however, focuses on more indirect approaches to inferring the objective from human feedback.
Modern reinforcement learning from human feedback (RLHF) originates from preference-based reinforcement learning (PbRL), independently introduced by Akrour et al. (2011) and Cheng et al. (2011). PbRL infers the objective from qualitative feedback, such as pairwise preferences between behaviors or between actions given states, instead of quantitative feedback in the form of numerical rewards. The term RLHF was later coined as an alternative (Askell et al., 2021; Ouyang et al., 2022; OpenAI, 2022), competing with the short-lived term reinforcement learning from human preferences (RLHP) (Menick et al., 2022), though initially referring to the same concept of learning behavior from relative feedback.
@@
Disentangling PbRL and RLHF is challenging due to their overlapping usage in the literature. For instance, Christiano et al. (2017) themselves are using the term PbRL, yet are often cited as a seminal reference for RLHF (Daniels-Koch & Freedman, 2022; Ouyang et al., 2022). This demonstrates that the terms are often used interchangeably. Practically, RLHF is often associated with reward modeling and deep RL, while PbRL is often linked to direct policy optimization in traditional RL settings. This is underlined by Jeon et al. (2020), who characterize PbRL as limited to direct policy learning from preferences. This is in contrast with other sources, however, who include reward learning within the scope of PbRL (Christiano et al., 2017; Wirth et al., 2017). Additionally, PbRL is rarely used in the language modeling domain, while RLHF is common in both language modeling and classical RL domains such as continuous control.
Despite overlapping and sometimes conflicting usage, in this work we view RLHF as a generalization of PbRL. While both involve human feedback to define RL objectives, PbRL primarily focuses on relative feedback, such as binary comparisons and rankings. RLHF not only includes these aspects but also extends to a wider range of feedback types (Metz et al., 2023; Yuan et al., 2024). Table˜1 gives an exemplary overview of our interpretation of these terms.
@@
We adopt the view that RLHF broadly encompasses all approaches using human feedback to define RL objectives, including both PbRL and SSRL. As the definitions and distinctions between these terms are not universally agreed upon, these distinctions are based on our interpretation of the current predominant usage of these terms in the literature.
@@ 1.3 Scope of the Survey
This section outlines the criteria guiding our selection of RLHF approaches we cover. We concentrate on methods where a learned reward model serves as the sole source of objective information, learned through interactive, online, scalable, and asynchronous human feedback. We describe each of these criteria in more detail below.
@@
We focus on approaches that learn a reward model from human feedback and then use this model to train a policy. Although it is possible to directly optimize a policy from human feedback (Wirth et al., 2017), thereby performing RLHF without reward learning, this approach was rarely practiced for a long time and has only recently gained renewed interest, especially in the domain of language model fine-tuning (see Section˜6.3). The decomposition into reward learning and policy training offers many conceptual and practical benefits. These include the direct applicability of supervised learning techniques for the reward model and the ability to evaluate the reward model in isolation. Additionally, the decomposition naturally leads to a form of semi-supervised learning, enabling the agent to use labeled episodes for reward model training while leveraging unlabeled episodes to refine its behavior and explore the environment.
@@
While many approaches include humans in the RL loop, we focus on approaches where human feedback is the only source of truth about the objective. This excludes approaches to reward shaping, feature engineering, and other forms of human guidance that are supplementary to a given objective.
@@
Our scope encompasses methods that collect human feedback interactively during the learning process, rather than relying solely on pre-collected demonstrations. This excludes pure imitation learning, learning from demonstration, and standalone inverse RL approaches. However, hybrid methods that combine inverse RL with interactive reward refinement fall within our scope, as discussed in Sections˜3.3 and 5.5.1.
@@
We consider methods that incorporate human feedback without requiring synchronous human-agent interaction. The agent must be capable of continuing its learning process while awaiting human input, and humans need not provide continuous supervision. This distinguishes RLHF from more direct methods of incorporating a human into the RL loop, and we believe that this is key for practicality and efficiency.
Beyond these methodological criteria, we primarily cover works from 2017 through 2024111We updated preprints to their peer-reviewed versions where applicable, leading to references outside this window., building upon the comprehensive treatment of earlier works by Wirth et al. (2017). We selectively revisit earlier foundational contributions that continue to inform current practice or have fundamentally shaped the field.
While RLHF has recently gained prominence through LLM applications, this survey adopts a broader perspective, examining RLHF across multiple domains with particular depth in control and robotics. We dedicate Appendix˜A to LLMs, addressing domain-specific challenges including large action spaces, distribution shift, and KL regularization. Many techniques from control domains transfer naturally to the LLM setting, which is frequently constrained to a single turn, though not all techniques are equally applicable. Given the rapidly evolving nature of LLM research, our coverage of techniques specific to this domain is less exhaustive than our treatment of control and robotics applications.
@@
This survey can be seen as the canonical continuation of Wirth et al. (2017), examining the evolution from PbRL to RLHF and the accompanying methodological advances. We provide both a thorough description of the basics and an in-depth discussion of current advances and trends. We refer to Appendix˜B for an overview of prior and related surveys and their relation to this work.